OpenStudy (anonymous):

solve cos x tan x - sin^2 x=0 for all real values of x??

OpenStudy (anonymous):

\[ \cancel{\cos x} \frac{\sin x}{\cancel \cos x} - \sin^2 x=0=>\sin x(1-\sin x)=0\] \[\sin x=0 ||\sin x=1\] \[x=0 || x= \frac{\pi}{2}\]

OpenStudy (anonymous):

Thank you!!